# Modeling temporal hypergraphs

### v2

@@ 1 Introduction
Graphs – collections of nodes and edges, where each edge connects two nodes – can represent networks of interconnected entities [1, 2]. Hypergraphs [3] generalize graphs by containing hyperedges that can connect any number of nodes. While having been suggested for network analysis decades ago [4], hypergraphs received a recent surge of interest [5, 6, 7, 8, 9, 10], in part due to the insight that crucial information gets lost when forcing hyperedges into collections of dyadic edges [11, 12]. Examples of complex systems that can be represented by hypergraphs are numerous and include networks of meetings or social events, emails, coauthoring and citation, co-offending, and chemical reactions [13, 14, 15, 16, 17, 18, 19]. Many of these examples give rise to temporal networks [20, 21], in this case, hypergraphs whose hyperedges are time-stamped or time-ordered [22].
@@ 2.1 Hypergraphs and temporal hypergraphs
A hypergraph is a pair $(\mathcal{I},H)$, where $\mathcal{I}=\{i_{1},\dots,i_{N}\}$ is a set of nodes (also denoted as actors in our main illustrating example) and $H\subseteq\mathcal{P}(\mathcal{I})$ is a set of hyperedges, that is, subsets of the node set. The size of a hyperedge $I\in H$ is the number of its nodes (i. e., $size(I)=|I|$) and the degree of a node $i\in\mathcal{I}$ in a hypergraph $(\mathcal{I},H)$ is the number of hyperedges containing $i$ (i. e., $deg(i)=|\{I\in H\colon i\in I\}|$). More generally, the degree of a set of nodes $I\subseteq\mathcal{I}$ is the number of hyperedges containing all nodes in $I$ (i. e., $deg(I)=|\{I^{\prime}\in H\colon I\subseteq I^{\prime}\}|$).
@@
For a given sequence of relational hyperevents $E$, that is, a temporal hypergraph, and a point in time $t$, let $E_{<t}=\{(t_{m},I_{m})\in E\colon t_{m}<t\}$ denote the history, also denoted as sequence of prior events, defined as the sub-sequence of all hyperedges that arrive strictly before $t$. We use this notation to define temporally changing degrees of node sets. For a set of nodes $I\subseteq\mathcal{I}$ and a point in time $t$, the prior degree of $I$ at $t$ is given by
$deg(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}\}|\kern 5.0pt.$
@@ 2.2 Data: Les Misérables
For illustration we use as a running example the network of actor co-appearance in chapters of Victor Hugo’s novel Les Misérables, compiled by Donald Knuth [52]. The data comprises 80 actors (co-)appearing in one or several of 288 chapters. Each of these chapters defines one hyperevent $e=(t,I)$, whose hyperedge $I$ contains the actors starring in that chapter and where the time stamps are counters for chapters $t\in\{1,\dots,288\}$ in the order in which they appear in the novel. To illustrate the use of exogenous node attributes, we augment Knuth’s Les Misérables data by a binary gender variable, $female(i)$, which is one if and only if actor $i$ is female. We believe that, using actors’ names, Knuth’s short description, and our own reading of the novel, we can derive this variable with good reliability. (We note that our analysis is given for illustrating the model and not for making any claims about the demography of the novel’s characters.) Among the 80 actors, there are 28 female (35%). For reproducibility, we provide the preprocessed data, augmented by the femalefemale variable, in the file jean_events.csv, linked from the eventnet Webpage  [^1] https://github.com/juergenlerner/eventnet/tree/master/data/les_miserables.
@@ 3.1 Model framework
Concretely, RHEM [44] specify the probability of a time-ordered sequence of hyperedges $E$, dependent on a vector of statistics $\bm{x}(\cdot,E_{<t_{m}})$ and an associated parameter vector $\bm{\beta}$, by
@@
where the conditional probability of the next hyperedge $I_{m}$, given the history of prior events, is specified by
@@ 3.2 Commonly used temporal hyperedge statistics
In the following, we introduce examples of frequently used temporal hyperedge statistics, building increasingly complex models, and briefly discuss selected findings on the Les Misérables data for illustration. A more detailed presentation and discussion of this analysis can be found in Appendix A.
@@
As it is common in real-world hypergraphs, node degrees are skewedly distributed in the Les Misérables data. For example, the most frequent actors are three of the novel’s main characters: Jean Valjean (JV, appearing in 113 chapters), Marius (MA, 77 chapters), and Cosette (CO, 55 chapters), largely exceeding the mean number of $9.1$ chapters per actor (compare Table 2).
@@
(The index set $\binom{I}{k}$ is the set of all subsets of size $k$, that is, $\binom{I}{k}=\{I^{\prime}\subseteq I\colon|I^{\prime}|=k\}$.) We point out that the previously discussed degsumdegsum statistic is identical to $subrep^{(1)}$. Figure 1 illustrates subset repetition of order one, two, and three. For example, subset repetition of order one takes higher values on the hyperedge $I^{\prime}=\{D,E,F\}$ while subset repetition of order three is higher for $I=\{A,B,C\}$, while both hyperedges have the identical value for subset repetition of order two.
@@
Figure 1: (Adapted from [12].) Illustrating subset repetition based on four prior events $e_{1},\dots,e_{4}$ on the hyperedges $I=\{A,B,C\}$ and $I^{\prime}=\{D,E,F\}$ at time $t>t_{4}$. For II it is $subrep^{(1)}(I^{\prime},E_{<t})=3$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$. For $I^{\prime}$ it is $subrep^{(1)}(I^{\prime},E_{<t})=6$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$.
Model 3 in Table 1 reveals that the Les Misérables data provides evidence for preferential attachment of subsets containing one, two, and three actors. Individual actors with a higher number of past events are selected into future events with a higher probability. Controlling for this individual popularity effect, pairs of actors that co-appear in more previous chapters are jointly selected into future chapters with a higher probability. Controlling for individual and pairwise popularity effects, triples of actors that co-appear in more previous chapters tend to co-appear in future chapters with a higher probability.
Table 1: Parameters of RHEM estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events. The numbering of models reflects that given in Appendix A, where we discuss further model specifications.
Model 3 Model 7 Model 9
@@
We can use any estimated RHEM also to compute the expected values of specifically chosen sets of nodes. For instance, we may use Model 1 in Table 3 in the appendix, specified only with preferential attachment of individual nodes, to check whether specific pairs of actors co-appear in chapters more or less frequently than suggested by their individual frequencies. We find that the most overrepresented dyad is JV-CO (expected to co-appear in 13 chapters vs. an observed number of 33 co-appearances), followed by EN-CR (expected 0.33; observed 18), EN-CM (expected 0.30, observed 15), and MA-CO (expected 4.7, observed 19). The dyad JV-MA, which is the one with the highest expected number of co-appearances (20 expected vs. 18 observed) is only on position 3,125 (among 3,160 dyads) when ranking dyads by their overrepresentation. So JV-MA is among the most underrepresented dyads. By Heider’s balance theory [58, 59, 60, 61], and labeling dyads with the differences of actually observed co-appearances minus the expected number, the triangle JV-MA-CO is an imbalanced triangle, where Jean Valjean and Marius are both positively linked to Cosette, but meet themselves less often than expected, given their individual frequencies.
Turning to overrepresentation or underrepresentation of filled triads, according to Model 1, we find that the most overrepresented triad is among three of the “friends of the ABC”, namely EN-CM-CR, which co-appear in 12 chapters (the maximum value over all triads) and are expected to co-appear in only 0.01 chapters. The triad formed by the three most frequent actors JV-MA-CO is expected to appear in 3.3 chapters but does appear in 6 chapters in the actual novel, placing this triad on position 164 (among 82,160 triads) when ranking triads by their overrepresentation according to Model 1. The assessment of the JV-MA-CO triad changes remarkably if we additionally control for a preferential attachment effect of actor pairs. According to Model 2 in Table 3 (Appendix), specified with subset repetition of order one and two, JV-MA-CO is the most underrepresented triad, with an expected number of 16 co-appearances but only 6 co-appearances in the actually observed data. This difference to the expectation of Model 1 can be explained by the finding that two of the three contained dyads, namely JV-CO and MA-CO, are strongly overrepresented (controlling for individual frequencies). Considering also previous dyadic co-appearances, the probability that all three, JV, MA, and CO, jointly appear in a chapter increases. This illustrates that if we want to assess the relative frequency of higher-order structures (such as filled triads), we should not only control for the prior degree of individual nodes but also for the prior occurrence of more complex sub-structures, such as dyads that are subsets of the respective triads. We can find a similar reasoning in [53] who argue that to test higher-order homophily in hypergraphs one has to control for lower-order homophily.
A variation of subset repetition is exact repetition, which captures a tendency to exactly repeat the participant list of prior events. This is a frequent pattern, for example, in email networks [39] (where senders repeatedly send messages to the same list of receivers) or in temporal hypergraphs representing meetings [44] (where regular meetings may have fixed attendee lists). We can test this effect with the statistic $repetition(I,E_{<t})$ counting the number of prior events on the hyperedge II
$repetition(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I=I_{m}\}|\kern 5.0pt.$ (6)
We find that the Les Misérables data provides evidence for exact repetition, in addition to subset repetition of order one, two, and three; see Table 4 in the appendix.
Triadic closure.
Triadic closure is the effect that two nodes $i_{1}$ and $i_{2}$ both had a prior contact with a common third node $i_{3}$, possibly in different past events, and then $i_{1}$ and $i_{2}$ meet in a future event. Triadic closure is a different effect than the repetition of filled triads. Indeed, even if all (hyper)edges have size two, there may still be triadic closure, but there cannot be any filled triads. We can test with RHEM whether a temporal hypergraph provides evidence for triadic closure, on top of what is explained by preferential attachment of individual nodes, dyads, or triads. For such tests we include a temporal hyperedge statistic $closure(I,E_{<t})$, defined by the sum:
$\sum_{\{i_{1},i_{2}\}\in\binom{I}{2}\atop i_{3}\in\mathcal{I}\colon i_{3}\neq i_{1},i_{2}}\min[deg(\{i_{1},i_{3}\},E_{<t}),deg(\{i_{2},i_{3}\},E_{<t})]\kern 5.0pt.$ (7)
For each dyad $\{i_{1},i_{2}\}$ in $I$ we look for “third nodes” $i_{3}$ (not necessarily in $I$) and then we weight the two-path $i_{1}$$i_{3}$$i_{2}$ by the minimum prior degree of $\{i_{1},i_{3}\}$ and $\{i_{2},i_{3}\}$. We then add up these weights over all dyads included in $I$ and all third nodes. Models in the second row of Table 3 in the appendix (also see Models 7 and 9 in Table 1) reveal that closure events are overrepresented and that this finding is robust if we control for repetition of individual nodes, dyads, and triads.
@@
The summation index runs over all unordered pairs $I^{\prime},I^{\prime\prime}$ of different subsets of size $k$ of the given hyperedge $I$. For each of these pairs we add the absolute difference between the degrees $deg(I^{\prime},E_{<t})$ and $deg(I^{\prime\prime},E_{<t})$. The resulting sum would be a measure of disassortativity as it gets larger when the prior degrees of the contained subsets are more different. We therefore take the negative sum to get a measure of degree assortativity. It is not necessary to add a constant in order to make the statistic positive, since any additive constant would cancel out in (3). We note that $assort^{(k)}(I,E_{<t})$ can assume non-zero values only if $|I|\geq k+1$, since there have to be more than one subset of size $k$ to get any pairs of such subsets.
In Model 7 in Table 1 we find that there is assortativity of individual nodes, so that actors tend to co-appear together with other actors having a similar prior degree (i. e., that appeared in a similar number of previous chapters). On top of that we find degree assortativity for pairs of nodes. This means that dyads tend to co-appear with other dyads having a similar number of previous co-appearances, controlling for assortativity of individual nodes.
@@
We can define statistics to assess first-order effects and homophily with respect to exogenous variables, such as age or gender. The definition of such statistics is very similar to subset repetition and assortativity of order one. The only change in the definition is that we replace the prior degree by a function giving the values of the respective variable. For example, if we have a binary variable $female(i)$, taking the value one if and only if $i$ is female, we can assess the first-order effect (testing whether females participate in events at a higher rate) via the statistic (which we also denote as femalefemale by a slight abuse of notation):
@@
$homophily(I)=-\sum_{\{i,i^{\prime}\}\in\binom{I}{2}}|female(i)-female(i^{\prime})|\kern 5.0pt.$ (10)
In the Les Misérables data (augmented by the gendergender variable as described above) we find that females are neither significantly overrepresented nor underrepresented in events (compared to their prevalence in the set of all 80 characters), but that there is homophily with respect to the “female” variable (see Model 9 in Table 1 and also compare Table 5 in the appendix). This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.

### v3

@@ 1 Introduction
Graphs – collections of nodes and edges, where each edge connects two nodes – can represent networks of interconnected entities [1, 2]. Hypergraphs [3] generalize graphs by containing hyperedges that can connect any number of nodes. While having been suggested for network analysis decades ago [4], hypergraphs received a recent surge of interest [5, 6, 7, 8, 9, 10], in part due to the insight that crucial information gets lost when forcing hyperedges into collections of dyadic edges [11, 12]. Examples of complex systems that can be represented by hypergraphs are numerous and include networks of meetings or social events, emails, coauthoring and citation, co-offending, and chemical reactions [13, 14, 15, 16, 17, 18, 19]. It has also been shown that higher-order interactions in evolutionary games can have implications on the emergence of collaboration [20, 21]. Many of these examples give rise to temporal networks [22, 23], in this case, hypergraphs whose hyperedges are time-stamped or time-ordered [24].
@@ 2.1 Hypergraphs and temporal hypergraphs
A hypergraph is a pair $(\mathcal{I},H)$, where $\mathcal{I}=\{i_{1},\dots,i_{N}\}$ is a set of nodes (also denoted as actors in our illustrating example) and $H\subseteq\mathcal{P}(\mathcal{I})$ is a set of hyperedges, that is, subsets of the node set. The size of a hyperedge $I\in H$ is the number of its nodes (i. e., $size(I)=|I|$) and the degree of a node $i\in\mathcal{I}$ in a hypergraph $(\mathcal{I},H)$ is the number of hyperedges containing $i$ (i. e., $deg(i)=|\{I\in H\colon i\in I\}|$). More generally, the degree of a set of nodes $I\subseteq\mathcal{I}$ is the number of hyperedges jointly containing all nodes in $I$ (i. e., $deg(I)=|\{I^{\prime}\in H\colon I\subseteq I^{\prime}\}|$).
@@
For a given sequence of relational hyperevents $E$, that is, a temporal hypergraph, and a point in time $t$, let $E_{<t}=\{(t_{m},I_{m})\in E\colon t_{m}<t\}$ denote the history, also denoted as sequence of prior events, defined as the sub-sequence of all hyperevents that arrive strictly before $t$. We use this notation to define temporally changing degrees of node sets. For a set of nodes $I\subseteq\mathcal{I}$ and a point in time $t$, the prior degree of $I$ at $t$ is defined as
$deg(I,E_{<t})=|\{(t_{m},I_{m})\in E_{<t}\colon I\subseteq I_{m}\}|\kern 5.0pt,$
that is, the number of all hyperevents arriving before $t$ and containing the entire set $I$.
@@ 2.2 Data: Les Misérables
For illustration we use as a running example the network of actor co-appearance in chapters of Victor Hugo’s novel Les Misérables, compiled by Donald Knuth [54]. The data comprises 80 actors co-appearing in one or several of 288 chapters. Each of these chapters defines one hyperevent $e=(t,I)$, whose hyperedge $I$ contains the actors starring in that chapter and where the time stamps are counters for chapters $t\in\{1,\dots,288\}$ in the order in which they appear in the novel. The number of actors per chapter, that is, the hyperedge size varies from one to ten, where chapters with more than four actors are rare; see Fig. 1.
To illustrate the use of exogenous node attributes, we augment Knuth’s Les Misérables data by a binary gender variable, $female(i)$, which is one if and only if actor $i$ is female. We believe that, using actors’ names, Knuth’s short description, and our own reading of the novel, we can derive this variable with good reliability. (We note that our analysis is given for illustrating the model and not for making any claims about the demography of the novel’s characters.) Among the 80 actors, there are 28 female (35%). For reproducibility, we provide the preprocessed data, augmented by the femalefemale variable, in the file jean_events.csv, linked from the eventnet Webpage  [^1] https://github.com/juergenlerner/eventnet/tree/master/data/les_miserables.
@@ 3.1 Model framework
Concretely, RHEM [46] specify the probability of a time-ordered sequence of hyperevents $E$, dependent on a vector of statistics $\bm{x}(\cdot,E_{<t_{m}})$ and an associated parameter vector $\bm{\beta}$, by
@@
that is, factorizing the probability of the entire sequence into the product of conditional probabilities of the next hyperedge, given the history of prior events. The conditional probability of the next hyperedge $I_{m}$, in turn, is specified by
@@ 3.2 Commonly used temporal hyperedge statistics
In the following, we introduce examples of frequently used temporal hyperedge statistics, building increasingly complex models, and briefly discuss selected findings on the Les Misérables data for illustration. We estimate model parameters by maximum likelihood estimation with case-control sampling, where we bound the number of selected non-events per event to $\binom{80}{3}=82,160$. This means that for events $(t_{m},I_{m})$ of size one, two, and three we take the entire set of subsets $\binom{\mathcal{I}}{|I_{m}|}$ – compare the index set in the denominator of (3) – and for larger events we bound the set of sampled non-events to the number of different triads of nodes. Sampling non-events and computing hyperedge statistics is done with the eventnet software [41]. Given the table of hyperedge statistics, parameter estimation is done with the coxph function in the R package survival [44, 45].
@@
As it is common in real-world hypergraphs, node degrees are skewedly distributed in the Les Misérables data. For example, the most frequent actors are three of the novel’s main characters: Jean Valjean (JV, appearing in 113 chapters), Marius (MA, 77 chapters), and Cosette (CO, 55 chapters), largely exceeding the mean number of $9.1$ chapters per actor.
@@
(The index set $\binom{I}{k}$ is the set of all subsets of size $k$, that is, $\binom{I}{k}=\{I^{\prime}\subseteq I\colon|I^{\prime}|=k\}$.) We point out that the previously discussed degsumdegsum statistic is identical to $subrep^{(1)}$. In general, subset repetition of order $k$ on a set of nodes $I$ is the sum of the prior degrees over all $k$-element subsets of $I$. Figure 2 illustrates subset repetition of order one, two, and three. For example, subset repetition of order one takes higher values on the hyperedge $I^{\prime}=\{D,E,F\}$, subset repetition of order three is higher for $I=\{A,B,C\}$, and both hyperedges have the identical value for subset repetition of order two.
@@
Figure 2: (Adapted from [12].) Illustrating subset repetition based on four prior events $e_{1},\dots,e_{4}$ on the hyperedges $I=\{A,B,C\}$ and $I^{\prime}=\{D,E,F\}$ at time $t>t_{4}$. For II it is $subrep^{(1)}(I^{\prime},E_{<t})=3$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$. For $I^{\prime}$ it is $subrep^{(1)}(I^{\prime},E_{<t})=6$, $subrep^{(2)}(I^{\prime},E_{<t})=3$, and $subrep^{(3)}(I^{\prime},E_{<t})=1$. Moreover, Event $e_{4}$ is an illustration of triadic closure (defined further below), as the event connecting Nodes $E$ and $F$, closes a two-path EEDDFF, which has been established in two previous events, $e_{2}$ and $e_{3}$.
Models reported in Table 1 reveal that the Les Misérables data provides evidence for preferential attachment of subsets containing one, two, and three actors. Individual actors with a higher number of past events are selected into future events with a higher probability. Controlling for this individual popularity effect, pairs of actors that co-appear in more previous chapters are jointly selected into future chapters with a higher probability. Controlling for individual and pairwise popularity effects, triples of actors that co-appear in more previous chapters tend to co-appear in future chapters with a higher probability.
Table 1: Parameters of RHEM estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events.
Model 1 Model 2 Model 3
@@
subrep(3) $0.218\;(0.024)^{***}$
AIC $3935.734$ $3626.535$ $3493.831$
$^{***}p<0.001$; $^{**}p<0.01$; $^{*}p<0.05$
We can use any estimated RHEM also to compute the expected values of specifically chosen sets of nodes. For instance, we may use Model 1 in Table 1, specified only with preferential attachment of individual nodes, to check whether specific pairs of actors co-appear in chapters more or less frequently than suggested by their individual frequencies. We find that the most overrepresented dyad is JV-CO (expected to co-appear in 13 chapters vs. an observed number of 33 co-appearances), followed by EN-CR (expected 0.33; observed 18), EN-CM (expected 0.30, observed 15), and MA-CO (expected 4.7, observed 19). The dyad JV-MA, which is the one with the highest expected number of co-appearances (20 expected vs. 18 observed) is only on position 3,125 (among 3,160 dyads) when ranking dyads by their overrepresentation. So JV-MA is among the most underrepresented dyads. By Heider’s balance theory [60, 61, 62, 63], and labeling dyads with the differences of actually observed co-appearances minus the expected number, the triangle JV-MA-CO is an imbalanced triangle, where Jean Valjean and Marius are both positively linked to Cosette, but meet themselves less often than expected, given their individual frequencies.
Turning to overrepresentation or underrepresentation of filled triads, according to Model 1, we find that the most overrepresented triad is among three of the “friends of the ABC”, namely EN-CM-CR, which co-appear in 12 chapters (the maximum value over all triads) and are expected to co-appear in only 0.01 chapters. The triad formed by the three most frequent actors JV-MA-CO is expected to appear in 3.3 chapters but does appear in 6 chapters in the actual novel, placing this triad on position 164 (among 82,160 triads) when ranking triads by their overrepresentation according to Model 1. The assessment of the JV-MA-CO triad changes remarkably if we additionally control for a preferential attachment effect of actor pairs. According to Model 2 in Table 1, specified with subset repetition of order one and two, JV-MA-CO is the most underrepresented triad, with an expected number of 16 co-appearances but only 6 co-appearances in the actually observed data. This difference to the expectation of Model 1 can be explained by the finding that two of the three contained dyads, namely JV-CO and MA-CO, are strongly overrepresented (controlling for individual frequencies). Considering also previous dyadic co-appearances, the probability that all three, JV, MA, and CO, jointly appear in a chapter increases. This illustrates that if we want to assess the relative frequency of higher-order structures (such as filled triads), we should not only control for the prior degree of individual nodes but also for the prior occurrence of more complex sub-structures, such as dyads that are subsets of the respective triads. We can find a similar reasoning in [55] who argue that to test higher-order homophily in hypergraphs one has to control for lower-order homophily.
Triadic closure.
Triadic closure is the effect that two nodes $i_{1}$ and $i_{2}$ both had a prior contact with a common third node $i_{3}$, possibly in different past events, and then $i_{1}$ and $i_{2}$ meet in a future event. Triadic closure is a different effect than the repetition of filled triads. Indeed, even if all (hyper)edges have size two, there may still be triadic closure, but there cannot be any filled triads. We can test with RHEM whether a temporal hypergraph provides evidence for triadic closure, on top of what is explained by preferential attachment of individual nodes, dyads, triads, or any other size of higher-order subsets. For such tests we include a temporal hyperedge statistic defined by:
$closure(I,E_{<t})=\sum_{\{i_{1},i_{2}\}\in\binom{I}{2}\atop i_{3}\in\mathcal{I}\colon i_{3}\neq i_{1},i_{2}}\min[deg(\{i_{1},i_{3}\},E_{<t}),deg(\{i_{2},i_{3}\},E_{<t})]\kern 5.0pt.$ (6)
For each dyad $\{i_{1},i_{2}\}$ in $I$ we look for “third nodes” $i_{3}$ (not necessarily in $I$) and then we weight the two-path $i_{1}$$i_{3}$$i_{2}$ by the minimum prior degree of $\{i_{1},i_{3}\}$ and $\{i_{2},i_{3}\}$. We then add up these weights over all dyads included in $I$ and all third nodes. Closure is illustrated in the stylized example shown in Fig. 2.
Models in Table 2 reveal that closure events are overrepresented in the Les Misérables data. We can also find specific chapters that are particularly well explained by triadic closure. Among the strongest is Chapter 4.15.2 in which Jean Valjean (JV) and Gavroche (GA) meet for the first time. Before that, both had prior encounters with various other common third actors, albeit in different chapters. The strongest “bundle” of prior two-paths connecting JV and GA is via Thénardier/Jondrette (TH/JD) and the gang of criminals (including MO, BB, BJ, GU) he collaborates with to extort money from JV in Chapter 3.8.20. Gavroche, on the other hand, assists in helping his (at the end unthankful) father TH to escape from prison in Chapter 4.6.3, together with MO, BB, BJ, GU, creating several two-paths between GA and JV before their encounter in Chapter 4.15.2.
Table 2: Parameters of RHEM estimated from the Les Misérables data. All models have been estimated with 288 events (chapters) and a cumulative number of $8,782,257$ selected non-events.
Model 4 Model 5 Model 6
[…]
subrep(2) $0.083\;(0.013)^{***}$ $0.102\;(0.013)^{***}$ $0.113\;(0.014)^{***}$
subrep(3) $0.126\;(0.032)^{***}$ $0.116\;(0.035)^{***}$ $0.121\;(0.035)^{***}$
closure $^{**} 0.008(0.003)0.008\;(0.003)^{**}$ $0.006\;(0.003)^{*}$
@@
The summation index runs over all unordered pairs $\{I^{\prime},I^{\prime\prime}\}$ of different subsets of size $k$ of the given hyperedge $I$. For each of these pairs we add the absolute difference between the degrees $deg(I^{\prime},E_{<t})$ and $deg(I^{\prime\prime},E_{<t})$. The resulting sum would be a measure of disassortativity as it gets larger when the prior degrees of the contained subsets are more different. We therefore take the negative sum to get a measure of degree assortativity. It is not necessary to add a constant in order to make the statistic positive, since any additive constant would cancel out in (3). We note that $assort^{(k)}(I,E_{<t})$ can assume non-zero values only if $|I|\geq k+1$, since there have to be more than one subset of size $k$ to get any pairs of such subsets.
In Model 5 in Table 2 we find that there is assortativity of individual nodes, so that actors tend to co-appear together with other actors having a similar prior degree (i. e., that appeared in a similar number of previous chapters). On top of that we find degree assortativity for pairs of nodes. This means that dyads tend to co-appear with other dyads having a similar number of previous co-appearances, controlling for assortativity of individual nodes.
@@
We can define statistics to assess first-order effects and homophily with respect to exogenous variables, such as age or gender. The definition of such statistics is very similar to subset repetition and assortativity of order one. The only change in the definition is that we replace the prior degree by a function giving the values of the respective variable. For example, if we have a binary variable $female(i)$, taking the value one if and only if $i$ is female, we can assess the first-order effect (testing whether females participate in events at a higher rate) via the statistic counting the number of females in a set of nodes $I$ (which we also denote as femalefemale by a slight abuse of notation):
@@
$homophily(I)=-\sum_{\{i,i^{\prime}\}\in\binom{I}{2}}|female(i)-female(i^{\prime})|\kern 5.0pt,$ (9)
counting the number of pairs in $I$ that have different gender.
In the Les Misérables data (augmented by the gendergender variable as described above) we find that females are neither significantly overrepresented nor underrepresented in events (compared to their prevalence in the set of all 80 characters), but that there is homophily with respect to the “female” variable; see Model 6 in Table 2. This means that actors typically co-appear with other actors of the same gender, more than what would be expected by chance alone.